Hello out there,I need some help with some tough problems,the first one is,Simplify: 9/k+8/k.The next one is,Find the quotient: 2x^2+7x+3/2x+1.The third one is,Find the quotient:(a^2+11a+30)divided by(2x+3).And the last one is,Find the quotient:(6x+9)divided by(2x+3),thanks to whoever can help.
Found 2 solutions by checkley71, babyb_:
Answer by checkley71(8403)

9/x+8/k=17/k answer.
--------------------------------------------------
(2x^2+7x+3)/(2x+1)
(2x+1)(x+3)/(2x+1) cancel out the (2x+1 terms.
x+3 answer.
--------------------------------------------------
(a^2+11a+30)/(2x+3)
(a+6)(a+5)/(2x+3) answer.
---------------------------------------------------
(6x+9)/(2x+3)
2(2x+3)/(2x+3) cancel out the 2x=3 terms.
2 answer.


Answer by babyb_(6)

(6x+9)/(2x+3)
do long division 2x+3 goes into 6x+9 3 times so the answer is 3 not 2.
2x times 3 = 6x and 3 times 3=9 so the answer is 3.